The longer leg of a right triangle is 6 less than 3 times the shorter leg. If the hypotenuse is 26cm., find the length of the two legs. Define the variables, set up your equation and solve.
Answer by mananth(16946)

let shorter leg be x
longer leg = 3x-6
hypotenuse = 26 cm
...
Pythagoras theorem
x^2+(3x-6)^2= 26^2
x^2+9x^2-36x+36=26^2
10x^2-36x+36-676=0
10x^2-36x-640=0
10x^2-100x+64x-640=0
10x(x-10)+64(x-10)=0
(10x+64)(x-10)=0
c= 10 which is positive
one leg is 10 cm
the other leg = 3x-6 = 30-6 =24cm
..
m.ananth@hotmail.ca